Find the total grazing area of a goat G, The animal is tied to a corner of a 40'x40' barn, by an 80' rope. One of the sides of the barn is extended by a fence. assume that the area is grass everywhere except inside of the barn. : Find the total grazing area of a goat G, The animal is tied to a corner of a 40'x40' barn, by an 80' rope. One of the sides of the barn is extended by a fence. assume that the area is grass everywhere except inside of the barn.
Answer by ankor@dixie-net.com(4538)

Find the total grazing area of a goat G, The animal is tied to a corner of a 40'x40' barn, by an 80' rope. One of the sides of the barn is extended by a fence. assume that the area is grass everywhere except inside of the barn.
;
The side that is extended by the fence will make a difference
:
If the fence is opposite the corner, where the rope is tied, then is will not effect it.
:
The larger area is 3/4 of circle with 80' radius
A = pi*80^2*.75
A = 15079.6 sq/ft
and
2 each 1/4 circles with a radius of 40'
A = pi*40^2*.25*2
A = 2513.3 sq/ft
:
Total: 15079.6 + 2513.3 = 17,593 sq/ft